An under water swimmer is at a depth of 12 m below the surface of water. A bird is at a height of 18 m from the surface of water, directly above his eyes. For the swimmer the bird appears to be at a distance ______ from the surface of water. (Refractive Index of water = 4/3 )
Answer : 24cm
1
Expert's answer
2019-07-08T10:26:35-0400

Using formula

"\frac{L'}{L}=n""n=4\/3; L=18 (m)""L'=n\cdot L= 24 (m)"